Soriano Company had net sales of $300,000 for the month (after returns and allowances of $1,500 and sales discounts of $3,250. Beginning inventory for the month was S60.000; purchases for the month were $175,000; and gross profit was 43%. What was the cost of goods sold for the month? Select one: O a. $304,750 b. $171,000 c. $235,000 d.$129,000 6 甸く※ 6 8 9. 0. 5
Answer #1

ANSWER is $171000

Gross Profit =43%

Hence Cost of goods sold =(1-43%)

Cost of Goods Sold=3000000*(1-.43) =$171000